OpenStudy (anonymous):

points a(1,2,3) and b(-1,0,6) a. find the midpoint of line segment AB b. find the length of line segment AB

sam (.sam.):

a(1,2,3) and b(-1,0,6) \[\Huge M(\frac{1+(-1)}{2},\frac{2+0}{2},\frac{3+6}{2})\]

OpenStudy (anonymous):

so it's 0, 1, 9/2?

sam (.sam.):

yes

OpenStudy (anonymous):

is the formula ( x1 + y1 / 2 , x2+y2 /2, x3+y3/2? :o

sam (.sam.):

\[\Huge M(\frac{x+x_1}{2},\frac{y+y_1}{2},\frac{z+z_1}{2})\]

OpenStudy (anonymous):

do you know how to find the length?

sam (.sam.):

\[L=\sqrt{(x-x_1)^2+(y-y_1)^2+(z-z_1)^2}\]

sam (.sam.):

\[L=\sqrt{17}\]

OpenStudy (anonymous):

ahhhh okay! thank you!!!